# Time-tronics: from temporal printed circuit board to quantum computer

### v1

@@
Let us assume that one atom occupies the wave-packet corresponding to the state $|1\rangle$ of a certain qubit, and another atom occupies the wave-packet corresponding to the state $|1\rangle$ of another qubit. Just before the moment when these two wave-packets meet, we can change the internal state of one or both atoms by activating the appropriate focused laser beams together with the broad laser beam, causing Raman transitions and altering the internal state of the atoms (Fig. 3b). Then, during the passage of the wave-packets, the atoms interact. Immediately after the passage, we use laser beams again to restore the atoms to their initial internal states, in which there are no interactions. If the interaction energy is properly chosen, a phase imprint of $e^{i\pi}$ occurs during the interaction. Such a phase imprint only occurs when both qubits are in the $|1\rangle$ states, thus implementing the controlled-Z gate. For all qubits whose wave-packets corresponding to the $|1\rangle$ states meet at the same moment, controlled-Z gate operations can be carried out in parallel. In subsequent periods of the oscillation of the optical lattice potential, a controlled-Z gate can be performed for other pairs of qubits whose wave-packets corresponding to the $|1\rangle$ states meet at different moments in time. To achieve high-fidelity controlled-Z gates, the entire gate operation needs to be divided into a few encounters of wave-packets (Fig. 1e), i.e., a few clock cycles of the quantum processor are necessary. The ability to perform two-qubit operations between all qubits makes a key parameter of quantum computing, i.e., quantum volume Cross et al. (2018), achieve very high values, especially when the Raman transfer error is reduced (Fig. 3c).
It is worth emphasizing that even with only a few focused laser beams, we can perform a network of quantum operations that are challenging to achieve in other experimental setups. A single beam focused in one location is sufficient to execute all necessary single-qubit operations, but in this case, the qubits must be associated with wave-packets that meet at the focus point of the beam. The addition of a second beam focused in another location allows the implementation of a controlled-Z gate between the nearest neighboring qubits that can be arranged in a 1D chain. If we introduce yet another focused beam, we gain the capability to perform controlled-Z gates between nearest neighboring qubits that can be arranged in a square network. A fourth beam enables controlled-Z gates between the nearest neighboring qubits arranged in a 3D cube, and so on.
@@
Periodically perturbed systems can give rise to crystalline structures in time, which, in the time domain, exhibit phenomena similar to those known in traditional spatial crystals Sacha (2020). In this study, we demonstrate that time crystalline structures possess significant potential for practical applications. In conventional spatial structures, various components of the system are spatially separated, and selective communication and interaction between distant parts of the system are preceded by their spatial transport. In time crystalline structures, the transport problem is automatically resolved because wave-packets associated with sites of time lattices evolve periodically in time, naturally encountering each other at different time instances. This opens up the possibility of creating any temporal printed circuit board capable of implementing a wide range of quantum devices for bosons or fermions. Similar to electronics, this gives rise to the field of time-tronics. What is a formidable challenge in spatial structures becomes trivial in temporal structures. As a concrete example, we describe the implementation of a universal quantum computer, in which the key problem of qubit transport Cirac and Zoller (2000); Bluvstein et al. (2022) for realizing two-qubit gates between any pair of qubits is automatically addressed.
@@ I Methods
Time crystalline structure
Consider a single atom confined within a 1D box potential periodically perturbed by an oscillating optical lattice potential (Fig. 1a). Using units for length, energy, and time, denoted as $1/k$, $\hbar^{2}k^{2}/m$, and $m/\hbar k^{2}$, respectively, where $k$ is the wave number of the laser beam creating the oscillating optical lattice potential, and $m$ is the mass of the atom, the system Hamiltonian is given by
@@
where $\lambda$ and $\omega$ represent the amplitude and frequency of the oscillating optical lattice potential, and $z$ ranges from 0 to the size of the box, $L=s\pi/2$, where $s$ is an integer. In the classical description, if the driving frequency is $s$ times larger than the frequency of the unperturbed motion of the atom, i.e., $\omega=s|p|\pi/L=2|p|$, where the resonant momentum $|p|\gg\sqrt{\lambda}$, the condition for the $s:1$ resonance is met Buchleitner et al. (2002). For a given optical lattice potential (i.e., for a selected wavelength of the laser beam, $2\pi/k$), the resonant number $s$ depends on the size of the 1D box potential. By increasing the size of the 1D box, we can achieve any value of $s$.
Before considering the quantum resonant behavior of the atom, let us commence with a classical description. We introduce a new pair of momentum and position variables, known as action-angle variables Buchleitner et al. (2002), $I=s|p|/2$ and $|\theta|=2z/s$ where $-\pi\leq\theta<\pi$. Switching to the moving frame, $\Theta=\theta-\omega t/s$, and neglecting rapidly oscillating terms, the effective Hamiltonian becomes $H_{\rm eff}=P^{2}/2+(\lambda s^{2}/32)\cos(s\Theta)$, where $P=I-\omega/s$ Buchleitner et al. (2002); Sacha (2020). For $s\gg 1$, $H_{\rm eff}$ describes an atom moving in a spatially periodic potential. In the quantum description, this results in solid-state-like behavior, which is manifested as the formation of energy bands and eigenstates in the form of Bloch waves Sacha (2020); Guo (2021).
Upon returning from the moving frame to the laboratory frame, this solid-state-like behavior is observed in the time domain Sacha (2020). Indeed, when fixing the position in space ($\theta=\rm const$) in the laboratory frame and examining how the probability of observing the atom at a chosen point changes in time, this replicates the crystalline structure observed versus $\Theta$ in the moving frame. This is due to the linear time transformation between the laboratory and moving frames, i.e., $\Theta=\theta-\omega t/s$. This concept has facilitated the exploration of various condensed matter phases in the time domain across different resonantly driven single-particle and many-body systems Sacha (2020); Hannaford and Sacha (2022).
@@
Control of the tunneling amplitudes $J_{ij}$
To selectively control atom transfer between different wave-packets, we can utilize Bragg scattering Müller et al. (2008). Let us consider two laser beams characterized by wave vectors $\mathbf{k}_{1}$ and $\mathbf{k}_{2}$ that we switch on for a short period of time $\tau$. The first beam is focused at the location where two wave-packets, $w_{i}(z,t)$ and $w_{j}(z,t)$, meet during evolution (Figs. 1a, 3a). The waist of the beam is larger than the width of the wave-packets but smaller than the distance between potential minima in (3). The second beam is broad and covers the entire 1D box. If the Bragg condition is satisfied, $\mathbf{k}_{1}-\mathbf{k}_{2}=2p\mathbf{e}_{z}$, an atom occupying the wave-packet $w_{j}$, moving with an average momentum $p\mathbf{e}_{z}$, is transferred to the other wave-packet $w_{i}$ which moves with an average momentum $-p\mathbf{e}_{z}$, and the reverse transfer is equally probable. These processes are described by an additional term in the Hamiltonian Müller et al. (2008)
@@
Control of the interaction strengths $U_{ij}$
Interactions between ultra-cold atoms are effectively short-range contact interactions described by a single parameter, the s-wave scattering length Pethick and Smith (2002). In the Hilbert subspace spanned by the wave-packets, $w_{i}(z,t)$, evolving along the periodic resonant trajectory, the interaction coefficients in the Hamiltonian (2) read Sacha (2020); Wang et al. (2021)
@@
with $g_{0}=2m\omega_{\perp}a/\hbar k$, where $a$ is the scattering length. Let us consider bosonic atoms where, choosing properly an external magnetic field in the vicinity of a Feshbach resonance Pethick and Smith (2002), we set the scattering length $a=0$. If we change the internal state of the atoms that occupy two wave-packets which are about to meet during the evolution along the periodic trajectory, they will interact because the scattering length $a$ is no longer zero. The change of the internal state can be performed by a focused laser beam which together with another broad beam realizes a Raman transition of the atoms between the internal states Levine et al. (2022) (Figs. 1c, 3b). Immediately after the wave-packets pass each other, the laser beams can be applied once again to restore the initial internal states of the atoms. The entire procedure leads to a time-dependent $g_{0}(t)$ in (5) which is non-zero only during the passing of the wave-packets. With an array of focused laser beams one can control and engineer all interaction coefficients $U_{ij}$ in the Hamiltonian (2), see Fig. 2b.
@@
For 39K atoms, which we are using as an example here, the s-wave scattering length can be set to zero if the atoms are initially prepared in the hyperfine state $|1,+1\rangle$ and the magnetic field is set to the zero crossing value of 350.5 G, close to the Feshbach resonance at 402.5 G D’Errico et al. (2007); Giergiel et al. (2020). However, if we transfer the atoms to the state $|1,-1\rangle$, which is 213 MHz above the $|1,+1\rangle$ state, using a Raman pulse, the corresponding scattering length becomes $-29$ Bohr radii, and the interaction coefficient (5) does not vanish. By varying the frequencies $\omega_{\perp}$ of the harmonic trap in the transverse directions (i.e., choosing different values of $g_{0}$), we can analyze how many meetings of the two wave-packets are needed to imprint a $\pi$ phase in the state in which each of the two atoms occupies one of the wave-packets. Simultaneously, we can determine the fidelity of the process, which is the probability that the atoms remain in their initially occupied wave-packets. The results of this analysis are presented in Fig. 1e. They correspond to $\omega_{\perp}/2\pi$ in the range between 50 kHz and 700 kHz.
Quantum computer
@@
Suppose that bosonic atoms are prepared in a Mott insulator state in a static optical lattice potential Pethick and Smith (2002), i.e., $\omega=0$ and $\lambda=7.5$ in (3). We assume that the number of atoms is equal to the number of lattice sites in the box potential, i.e., $s/2$ (Fig. 3), i.e., there is one atom in each potential well. After preparing the Mott insulator phase, we adjust the s-wave scattering length to zero using the Feshbach resonance as described in the preceding paragraph. We also kick the atoms so that they start moving in the same direction with momentum $p$ and activate resonant oscillations of the optical lattice with frequency $\omega=2|p|=60$ and amplitude $\lambda=30$. As a result, we have $s/2$ wave-packets occupied by single atoms, which start moving with average momentum $p$ along the resonant trajectory corresponding to the $s:1$ resonance. At the same time, the remaining $s/2$ resonant wave-packets which are not occupied by atoms start moving with average momentum $-p$. Now we can associate the states of $s/2$ qubits with wave-packets. We assign one occupied wave-packet and one unoccupied wave-packet to each qubit. For example, occupied wave-packets will be called states $|1\rangle$, and unoccupied wave-packets will be called states $|0\rangle$ of the qubits. With this convention, initially, we have all qubits prepared in the state $|1\rangle$.
@@
When the wave-packets corresponding to the states $|0\rangle$ and $|1\rangle$ of the same qubit meet during evolution along the resonant trajectory, Bragg scattering allows for $\sigma_{x}$ or $\sigma_{y}$ single-qubit operations depending on the relative phase $\phi$ between the broad laser beam and the beam focused at the meeting point of the wave packets (Fig. 3a). Figure 1d presents the fidelity of the single-qubit gates for the system of 39K atoms which we consider as an example here. A single focused laser beam (together with a broad beam) is sufficient to perform single-qubit operations for all qubits if the qubit states are assigned so that the wave-packets corresponding to $|0\rangle_{i}$ and $|1\rangle_{i}$ pass each other at different moments of time but at the same location for all $i$, i.e., all qubits. On the other hand, if multiple focused laser beams are available, the appropriate assignment allows for parallel single-qubit operations.
@@
To implement a controlled-Z gate for the $i$-th and $j$-th qubits, we need to wait until the wave-packets corresponding to the states $|1\rangle_{i}$ and $|1\rangle_{j}$ of those qubits meet during evolution along the resonant trajectory. Just before their meeting, we briefly turn on the broad laser beam and the laser beam focused at the location of the wave-packets, which induce Raman transfer to the internal states of the atoms where the atoms interact Levine et al. (2022) (Fig. 3b). After the wave-packets pass each other, the Raman laser beams are briefly turned on again, causing the atoms to return to their initial non-interacting internal states. If the interaction strength and the duration of the interaction are properly chosen, after the wave-packets pass each other, the state $|1\rangle_{i}|1\rangle_{j}$ acquires a phase $e^{i\pi}$ and a controlled-Z gate is realized. To minimize coupling of the wave-packets with other states during the interaction, i.e., to achieve high fidelity of the gate, it is necessary to divide the gate implementation into multiple encounters of the wave-packets (Fig. 1e).
@@
For the 39K atoms under consideration and for the fidelity, $F_{R}=0.999933$, achieved in the Raman transfer in Ref. Bluvstein et al. (2022), the optimal fidelity of the controlled-Z gate is 0.987 which requires 14 cycles of the resonant trajectory (Fig. 1e). The fidelity was determined by multiplying the fidelity obtained assuming perfect Raman transfers (black curve in Fig. 1e) by $F_{R}^{8n}$, where the $8n$ comes from the fact that for each of $n=14$ cycles two interactions take place where two atoms are transferred twice by Raman pulses. Higher fidelity of the gate can be achieved either by improving $F_{R}$ or by increasing the frequency of the optical lattice oscillation $\omega$. A higher $\omega$ allows for a greater oscillation amplitude $\lambda$, thereby enabling the use of stronger interactions without undesired couplings to other states of the system. Stronger interactions facilitate imprinting the $\pi$ phase onto states with high fidelity in fewer cycles of the resonant trajectory, thereby reducing the impact of imperfect Raman transfers.
@@
A significant advantage of the proposed quantum computer is the ability to perform two-qubit gates between any pair of qubits. This is quantitatively illustrated by the quantum volume presented in Fig. 3c. In the example we consider here and for the Raman transfer fidelity achieved in Ref. Bluvstein et al. (2022), $F_{R}=0.999933$, and for 49 cycle single-qubit gates (cf. Fig. 1d), a system of 7 qubits is capable of executing universal quantum operations.

### v2

@@
Let us assume that one atom occupies the wave-packet corresponding to the state $|1\rangle$ of a certain qubit, and another atom occupies the wave-packet corresponding to the state $|1\rangle$ of another qubit. Just before the moment when these two wave-packets meet, we can change the internal state of the atoms by activating the appropriate focused laser beams together with the broad laser beam, causing Raman transitions and altering the internal state of the atoms (Fig. 3b). Then, during the passage of the wave-packets, the atoms interact. Immediately after the passage, we use laser beams again to restore the atoms to their initial internal states, in which there are no interactions. If the interaction energy is properly chosen, a phase imprint of $e^{i\pi}$ occurs during the interaction. Such a phase imprint only occurs when both qubits are in the $|1\rangle$ states, thus implementing the controlled-Z gate. To achieve high-fidelity controlled-Z gates, the entire gate operation needs to be divided into a few encounters of wave-packets (Fig. 1e-1f), i.e., a few clock cycles of the quantum processor are necessary. All pairs of wave-packets meet each other during a single period, $sT$, of the resonant trajectory. Since the controlled-Z gate implementation does not alter the wave-packet populations of atoms, it can be executed simultaneously even for all possible qubit pairs, i.e., for $s/2(s/2-1)$ pairs, significantly reducing quantum computation time. The ability to perform two-qubit operations between all qubits makes a key parameter of quantum computing, i.e., the quantum volume [47], achieve very high values, especially when the Raman transfer error is reduced (Fig. 3c).
We emphasize that even with only a few focused laser beams, we can perform a network of quantum operations that are challenging to achieve in other experimental setups. A single beam focused in one location is sufficient to execute all necessary single-qubit operations, but in this case, the qubits must be associated with wave-packets that meet at the focus point of the beam. The addition of a second beam focused in another location allows the implementation of a controlled-Z gate between the nearest neighboring qubits that can be arranged in a 1D chain. If we introduce yet another focused beam, we gain the capability to perform controlled-Z gates between nearest neighboring qubits that can be arranged in a square network. A fourth beam enables controlled-Z gates between the nearest neighboring qubits arranged in a 3D cube, and so on.
@@
Periodically perturbed systems can give rise to crystalline structures in time, which, in the time domain, exhibit phenomena similar to those known in traditional spatial crystals [1]. In this study, we demonstrate that time crystalline structures possess significant potential for practical applications. In conventional spatial structures, various components of the system are spatially separated, and selective communication and interaction between distant parts of the system are preceded by their spatial transport. In time crystalline structures, the transport problem is automatically resolved because wave-packets associated with sites of time lattices evolve periodically in time, naturally encountering each other at different time instances. This opens up the possibility of creating any temporal printed circuit board capable of implementing a wide range of quantum devices for bosons or fermions. Similar to electronics, this gives rise to the field of time-tronics. What is a formidable challenge in spatial structures becomes trivial in temporal structures. As a concrete example, we describe the implementation of a universal quantum computer, in which the key problem of qubit transport [49, 43] for realizing two-qubit gates between any pair of qubits is automatically addressed.
@@ I Methods
Time crystalline structure. Consider a single atom confined within a 1D box potential periodically perturbed by an oscillating optical lattice potential (Fig. 1a). Using units of length, energy, and time, denoted by $1/k$, $\hbar^{2}k^{2}/m$, and $m/\hbar k^{2}$, respectively, where $k$ is the wave number of the laser beam creating the oscillating optical lattice potential, and $m$ is the mass of the atom, the system Hamiltonian is given by
@@
where $\lambda$ and $\omega$ represent the amplitude and frequency of the oscillating optical lattice potential, and $z$ ranges from 0 to the size of the box, $L=s\pi/2$, where $s$ is an integer. In the classical description, if the driving frequency is $s$ times larger than the frequency of the unperturbed motion of the atom, i.e., $\omega=s|p|\pi/L=2|p|$, where the resonant momentum $|p|\gg\sqrt{\lambda}$, the condition for the $s:1$ resonance is met [29]. For a given optical lattice potential (i.e., for a selected wavelength of the laser beam, $2\pi/k$), the resonant number $s$ depends on the size of the 1D box potential. By changing the size of the 1D box, we can achieve any value of $s$.
Before considering the quantum resonant behavior of the atom, let us commence with a classical description. We introduce a new pair of momentum and position variables, known as action-angle variables [29], $I=s|p|/2$ and $|\theta|=2z/s$ where $-\pi\leq\theta<\pi$. Switching to the moving frame, $\Theta=\theta-\omega t/s$, and neglecting rapidly oscillating terms, the effective Hamiltonian becomes $H_{\rm eff}=P^{2}/2+(\lambda s^{2}/32)\cos(s\Theta)$, where $P=I-\omega/s$ [29, 1]. For $s\gg 1$, $H_{\rm eff}$ describes an atom moving in a spatially periodic potential. In the quantum description, this results in solid-state-like behavior, which is manifested as the formation of energy bands and eigenstates in the form of Bloch waves [1, 30].
Upon returning from the moving frame to the laboratory frame, this solid-state-like behavior is observed in the time domain [1]. Indeed, when fixing the position in space ($\theta=\rm const$) in the laboratory frame and examining how the probability of observing the atom at a chosen point changes in time, this replicates the crystalline structure observed versus $\Theta$ in the moving frame. This is due to the linear time transformation between the laboratory and moving frames, i.e., $\Theta=\theta-\omega t/s$. This concept has facilitated the exploration of various condensed matter phases in the time domain across different resonantly driven single-particle and many-body systems [1, 3].
@@
Control of the tunneling amplitudes $J_{ij}$. To selectively control atom transfer between different wave-packets, we can utilize Bragg scattering [40]. Let us consider two laser beams characterized by wave vectors $\mathbf{k}_{1}$ and $\mathbf{k}_{2}$ that we switch on for a short period of time $\tau$. The first beam is focused at the location where two wave-packets, $w_{i}(z,t)$ and $w_{j}(z,t)$, meet during evolution (Figs. 1b, 3a). The waist of the beam is larger than the width of the wave-packets but smaller than the distance between potential minima in (3). The second beam is broad and covers the entire 1D box. If the Bragg condition is satisfied, $\mathbf{k}_{1}-\mathbf{k}_{2}=2p\mathbf{e}_{z}$, an atom occupying the wave-packet $w_{j}$, moving with an average momentum $p\mathbf{e}_{z}$, is transferred to the other wave-packet $w_{i}$, which moves with an average momentum $-p\mathbf{e}_{z}$, and the reverse transfer is equally probable. These processes are described by an additional term in the Hamiltonian [40]
@@
Control of the interaction strengths $U_{ij}$. Interactions between ultra-cold atoms are effectively short-range contact interactions described by a single parameter, the s-wave scattering length [44]. In the Hilbert subspace spanned by the wave-packets, $w_{i}(z,t)$, evolving along the periodic resonant trajectory, the interaction coefficients in the Hamiltonian (2) read [1, 51]
@@
with $g_{0}=2m\omega_{\perp}a/\hbar k$, where $a$ is the scattering length. Let us consider bosonic atoms where, choosing properly an external magnetic field in the vicinity of a Feshbach resonance [44], we set the scattering length $a=0$. If we change the internal state of the atoms that occupy two wave-packets which are about to meet during the evolution along the periodic trajectory, they will interact because the scattering length $a$ is no longer zero. The change of internal state can be performed by a focused laser beam which together with another broad beam realizes a Raman transition between the internal states [42] (Figs. 1c, 3b). Immediately after the wave-packets pass each other, the laser beams can be applied once again to restore the initial internal states of the atoms. The entire procedure leads to a time-dependent $g_{0}(t)$ in (5) which is non-zero only during the passing of the wave-packets. With an array of focused laser beams one can control and engineer all interaction coefficients $U_{ij}$ in the Hamiltonian (2), see Fig. 2b.
@@
For 39K atoms, which we use as an example here, the s-wave scattering length can be set to zero if the atoms are initially prepared in the hyperfine state $|1,+1\rangle$ and the magnetic field is set to the zero crossing value of 350.5 G, in the vicinity of the Feshbach resonance at 402.5 G [52, 53]. However, if we transfer the atoms to the state $|1,-1\rangle$, which is 213 MHz above the $|1,+1\rangle$ state, using a Raman pulse, the corresponding scattering length becomes $-29$ Bohr radii, and the interaction coefficient (5) does not vanish. By varying the frequencies $\omega_{\perp}$ of the harmonic trap in the transverse directions (i.e., choosing different values of $g_{0}$), we can analyze how many meetings of the two wave-packets are needed to imprint a $\pi$ phase in the state in which each of the two atoms occupies one of the wave-packets. Simultaneously, we can determine the fidelity of the process, which is the probability that the atoms remain in their initially occupied wave-packets. The results of this analysis are presented in Fig. 1e-1f. They correspond to $\omega_{\perp}/2\pi$ in the range between 50 kHz and 700 kHz.
@@
Quantum computer. Suppose that bosonic atoms are prepared in a Mott insulator state in a static optical lattice potential [44], i.e., $\omega=0$ and $\lambda=7.5$ in (3). We assume that the number of atoms is equal to the number of lattice sites in the box potential, $s/2$ (Fig. 3), i.e., there is one atom in each potential well. After preparing the Mott insulator phase, we adjust the s-wave scattering length to zero using a Feshbach resonance as described in the preceding paragraph. We also set the atoms moving in the same direction with momentum $p$ using the counter-diabatic driving process (see Supplemental Material) that satisfies the $s:1$ resonance condition and activate resonant oscillations of the optical lattice with frequency $\omega=2|p|=60$ and amplitude $\lambda=30$. As a result, we have $s/2$ wave-packets occupied by single atoms, which start moving with average momentum $p$ along the resonant trajectory corresponding to the $s:1$ resonance. At the same time, the remaining $s/2$ resonant wave-packets which are not occupied by atoms start moving with average momentum $-p$. Now we can associate the states of $s/2$ qubits with wave-packets. We assign one occupied wave-packet and one unoccupied wave-packet to each qubit. For example, occupied wave-packets will be called states $|1\rangle$, and unoccupied wave-packets will be called states $|0\rangle$ of the qubits. With this convention, initially, we have all qubits prepared in the state $|1\rangle$.
@@
When the wave-packets corresponding to the states $|0\rangle$ and $|1\rangle$ of the same qubit meet during evolution along the resonant trajectory, Bragg scattering allows for $\sigma_{x}$ or $\sigma_{y}$ single-qubit operations depending on the relative phase $\phi$ between the broad laser beam and the beam focused at the meeting point of the wave-packets (Fig. 3a). Figure 1d presents the fidelity of the single-qubit gates for the system of 39K atoms which we consider as an example here. A single focused laser beam (together with a broad beam) is sufficient to perform single-qubit operations for all qubits if the qubit states are assigned so that the wave-packets corresponding to $|0\rangle_{i}$ and $|1\rangle_{i}$ pass each other at different moments of time but at the same location for all $i$, i.e., all qubits. On the other hand, if multiple focused laser beams are available, the appropriate assignment allows for parallel single-qubit operations.
@@
To implement a controlled-Z gate for the $i$-th and $j$-th qubits, we need to wait until the wave-packets corresponding to the states $|1\rangle_{i}$ and $|1\rangle_{j}$ of those qubits meet during evolution along the resonant trajectory. Just before their meeting, we briefly turn on the broad laser beam and the laser beam focused at the location of the wave-packets, which induce Raman transfer to the internal states of the atoms where the atoms interact [42] (Fig. 3b). After the wave-packets pass each other, the Raman laser beams are briefly turned on again, causing the atoms to return to their initial non-interacting internal states. If the interaction strength and the duration of the interaction are properly chosen, after the wave-packets pass each other, the state $|1\rangle_{i}|1\rangle_{j}$ acquires a phase $e^{i\pi}$ and a controlled-Z gate is realized. To minimize coupling of the wave-packets with other states during the interaction, i.e., to achieve high fidelity of the gate, it is necessary to divide the gate implementation into multiple encounters of the wave-packets (Fig. 1e-1f).
@@
For the 39K atoms under consideration and for the fidelity, $F_{R}=0.999933$, achieved in the Raman transfer in Ref. [43], the optimal fidelity of the controlled-Z gate is 0.993 which requires 7 cycles of the resonant trajectory (Fig. 1f). The fidelity was determined by multiplying the fidelity obtained assuming perfect Raman transfers (black curve in Fig. 1f) by $F_{R}^{8n}$, where the $8n$ comes from the fact that for each of $n=7$ cycles two interactions take place where two atoms are transferred twice by Raman pulses. Higher fidelity of the gate can be achieved either by improving $F_{R}$ or by increasing the frequency of the optical lattice oscillation $\omega$. A higher $\omega$ allows for a greater oscillation amplitude $\lambda$, thereby enabling the use of stronger interactions without undesired couplings to other states of the system. Stronger interactions facilitate imprinting the $\pi$ phase onto states with high fidelity in fewer cycles of the resonant trajectory, thereby reducing the impact of imperfect Raman transfers.
@@
A significant advantage of the proposed quantum computer is the ability to perform two-qubit gates between any pair of qubits. This is quantitatively illustrated by the quantum volume presented in Fig. 3c. In the example we consider here and for the Raman transfer fidelity achieved in Ref. [43], $F_{R}=0.999933$, and for 49 cycle single-qubit gates (cf. Fig. 1d), a system of 7 qubits is capable of executing universal quantum operations. In this case, achieving the $14:1$ resonance is necessary, and for 39K atoms, any combination of the $\sigma_{x}$ and $\sigma_{y}$ operations takes 126 ms, while the controlled-Z gate operation takes 36 ms. We note that during these periods all single-qubit operations and controlled-Z gate operations, even for all possible qubit pairs, can be executed simultaneously, significantly reducing quantum computation time.